# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v2

@@ Abstract
We consider the unitary quantum dynamics of a Bose-Einstein condensates at absolute zero, and demonstrate that an analogue gravity model in such a setting must take into account the backreaction of quasiparticle excitations onto the condensate background. This in turn requires that one expands to second order in perturbations, and takes the nonlinearity of the theory into account. It is shown that this leads to significant modifications of the standard linearized analogue gravity paradigm à la Unruh. In particular, to obtain a fully Lorentz-covariant equation in curved spacetime for second-order perturbations, we demonstrate that it is necessary to introduce, to leading order in powers of the formal expansion parameter $N^{-1/2}$ (where $N$ is total particle number), a quantum-fluctuation-renormalized spacetime metric which substantially differs from the Unruh metric and, to subleading order, two emergent vector fields. Both the renormalized metric and the vector fields then keep track of the backreaction of the quasiparticles onto the condensate up to the order in powers of $N^{-1/2}$ considered. Finally, we apply our formalism to an analogue-cosmological Friedmann-Lemaître-Robertson-Walker metric and establish its renormalized form due to the quantum many-body backreaction exerted by the excitation cloud.
@@ I Introduction
It is of major importance in what follows that for a proper treatment of backreaction problems in quantum systems such as BECs, as was first pointed out in [30], it is necessary to maintain particle-number conservation, cf. Refs. [31, 32, 33, 34], which are obtaining variants of number-conserving formulations of Bogoliubov theory which go beyond the usual symmetry-breaking, traditional Bogoliubov approach. To this end, in [30] the backreaction problem was defined by employing an expansion of the field operator in powers of $N^{1/2}$, where $N$ is the total number of particles (here either atoms or molecules) in the system. Then the backreaction of the quantum fluctuations can be elegantly treated as the subleading correction to the leading order in $N^{1/2}$ Gross-Pitaevskiǐ equation (GPE) [equivalent to a nonrelativistic variant of $\phi^{4}$ theory], incorporated by the modified coupled dynamical equation of the condensate and the non-condensate part. A particular lesson to be drawn from the analysis of [30] is that in a fully number-conserving description of the backreaction problem, the usual assumption of taking the expectation value of the quantized energy-momentum tensor, $\hat{T}_{\mu\nu}$, in a suitable quantum state, cf. Ref. [35], does not necessarily reproduce the correct, that is observable physical backreaction force.
@@
We will argue that the above procedure engenders a significant departure from the established formulation of the fluid perturbation equations as those of a scalar field in curved spacetime and that it is not much less straightforward to conclude that such a description can be envisaged if both backreaction and nonlinearity are taken into account. We will indeed demonstrate that for a Lorentz-invariant form of the nonlinear perturbation equation for a scalar field to exist, the latter will acquire, in subleading order, a mass term and two “emergent” vector fields. To leading order in the large $N$ expansion, the effective curved background metric acquires corrections stemming from the exchange of particles between condensate and excitation cloud. As our analysis in this paper is based on the Bogoliubov approximation, i.e., the number of particles forming the condensate must be much larger than that of forming the depleted cloud, the backreaction corrections to the standard GPE are by necessity small (for the Bogoliubov theory to apply the gas must be dilute), and consequently also the above mentioned differences from the standard linearized analogue gravity procedure (which we dub in what follows the Unruh paradigm) must be small. However, due to the high tunability of quantum gases, in particular with respect to the two-body coupling strength, by Feshbach resonances (cf., e.g., Refs. [43, 44]), one can potentially observe such deviations from the conventional analogue gravity paradigm in a highly controlled manner in cold quantum gas experiments.
The paper is organized as follows: In Section. II, we explain the formulation of the number-conserving approach we follow throughout the paper. Section III consists of the description of identifying the Lorentz invariant dynamical equation for the perturbed field. Section IV describes the quantum nonlinear corrections to the lowest order semiclassical background metric and Section V elaborates on these corrections for a Friedmann-Lemaître-Robertson-Walker metric. In Section VI we conclude, and an Appendix contains the explicit evaluation of the drain term driving condensate loss for a Bose gas in a hard-walled box trap which is quenched from zero to finite interaction coupling.
@@ II.1 Governing field equations
$i\partial_{t}\Psi=\left(-\frac{1}{2}\nabla^{2}+V+U|\Psi|^{2}\right)\Psi,$ (1)
Here, $V$ is the external (trapping) potential, and $U>0$ is the interaction coupling from $s$-wave scattering between the gas particles. The theory described by (1) is invariant under global $U(1)$ transformations, which ensures the conservation of the total particle number $N$ according to the Noether theorem. The conservation law $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$ thus holds, where
@@
Here, the scaling behavior of each components taking the limit of very large $N$ while keeping $UN=\mbox{constant}$, is given by [30]
@@
The above formal expansion (3) relies on the assumption that the number of non-condensed particles, $\delta N$, is much less than that of those forming the condensate part, and thus posits that $N_{0}\lesssim N$, also cf. Ref. [46] for the proof that in the mathematical limit $N\rightarrow\infty$ with $UN$ held fixed, the GPE (5) becomes exact. Here, $\chi$ is the fluctuation field associated with the non-condensed particles, and $\phi_{2}$ encompasses the correction of the condensate wave function due to the backreaction of the field $\chi$ onto the condensate. An expansion respecting the $N$ hierarchy is necessary to properly identify the fluctuation operator and the correction to the leading order Gross-Pitaevskiǐ orbital, see for details below.
@@
$i\frac{\partial}{\partial t}\phi_{0}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{0}|^{2}\right]\phi_{0}.$ (5)
@@
$i\frac{\partial}{\partial t}\chi=\left(-\frac{1}{2}\nabla^{2}+V+2U\rho_{0}\right)\chi+U\phi_{\rm 0}^{2}\chi^{*},$ (6)
with $\rho_{0}=|\phi_{0}|^{2}$. The above equation and the corresponding Hermitian conjugate equation are the modified versions of the Bogoliubov–de Gennes equations in the non-number-conserving standard Bogoliubov approach. When we consider the next-order corrections to the condensate wave function due to the backreaction of the non-condensate part of the system, as was done in [33, 34], then it can be shown that the $\mathcal{O}(N^{0})$ corrections to the leading order GPE are identically zero [33]. However in the next order ${\mathcal{O}}(N^{-1/2})$, the standard GPE is modified due to the correction from the time-dependent backreaction of the noncondensate particles, as described by the fluctuation operator. To provide a description of the corresponding mechanism, we will follow [30, 36] and define a quantum-fluctuation-corrected condensate wave function
@@
$i\partial_{t}\phi_{c}=\Big{(}-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U|\chi|^{2}\Big{)}\phi_{c}+U\chi^{2}\phi_{c}^{*}.$ (8)
This equation describes the coupled dynamics of condensed and noncondensed parts taking the dynamical corrections due to the backreaction into account. We note that the corrected order parameter $\phi_{c}$ has a dominant ${\cal O}(N^{1/2})$ term plus a subdominant correction of order $N^{-1/2}$, and for a proper formulation of the backreaction, it is essential to take this correction into account.
@@
$\langle\hat{\chi}({\bm{x}},t))\rangle\coloneqq 0\,\,\,\forall\,{\bm{x}},t\geq 0.$ (10)
This quantization choice, that is to regard $\hat{\chi}$ as the fundamental bosonic quantum field instead of $\hat{\Psi}$, does not maintain number conservation on the level of the full field operator of the bosonic particles: Due to $\langle\hat{\Psi}\rangle=\phi_{c}$, where however now $\phi_{c}$ is the backreaction-renormalized condensate mode in Eq. (7), the U(1) symmetry is still broken on the quantum level, as in the original Bogoliubov theory [27]. This choice, to promote $\chi$ to be the “fundamental” quantum field, however yields a set of equations amenable to a concrete solution the backreaction problem  [^2] We note in this connection that we were unable to locate an explicit backreaction calculation following the formally rigorous and in this regard very appealing approach of Ref. [33], which maintains $\langle\hat{\Psi}\rangle=0$ for any fixed $N$ state.. Furthermore, and most importantly in our context, it has been shown in [36] that with this vacuum choice for the quantized field $\hat{\chi}$, number conservation for condensate plus excitations is maintained, see Eqs. (18) and (19) below, and up to the order in $N^{-1/2}$ considered. The initial state for the Heisenberg picture evolution is then chosen as the noninteracting $U=0$ which has zero depletion $\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle=0\,\,(t=0)$ and $\phi_{2}=0$ (zero backreaction) [36] and can be set up in experiment in a well controlled manner. At $t=0$, we quench to a finite value of $U$ and follow the subsequent evolution.
@@
$i\partial_{t}\phi_{c}=\left[-\frac{1}{2}\nabla^{2}+V+U|\phi_{\rm c}|^{2}+2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle\right]\phi_{c}+U\langle\hat{\chi}^{2}\rangle\phi_{c}^{*}.$ (11)
@@ II.2 Quantum backreaction
$\rho$ $=$ $|\phi_{c}({\bm{x}}|^{2}+\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+{\mathcal{O}}(N^{-1/2}),$ (14)
${\bm{J}}$ $=$ $\frac{1}{2i}\Big{(}\phi^{*}_{c}\nabla\phi_{c}({\bm{x}}-\nabla\phi^{*}_{c}\Big{)}$ (15)
@@
We then have $\rho=\rho_{c}+\rho_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$ and $J=J_{c}+J_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$.
Neither condensate nor depletion are separately conserved, even if only up to the order ${\mathcal{O}}(N^{0})$ we consider. In particular, we have for condensate and depletion cloud, respectively [36, 37],
$\partial_{t}\rho_{c}+\nabla\cdot{\bm{J}}_{c}$ $=$ $iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}$ (18)
$\partial_{t}\rho_{\hat{\chi}}+\nabla\cdot{\bm{J}}_{\hat{\chi}}$ $=$ $-iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}.$ (19)
We conclude that the theory is indeed conserving up to order ${\mathcal{O}}(N^{0})$ for density and current. This defines the extent to which our approach is number-conserving, as required. The number of particles in condensate and in the depletion cloud of quantum excitations are not separately conserved but only their respective sums, and thus the total density and current. The separate conservation equations for condensate and excitations acquire drain (source) terms which stem from backreaction.
@@ III Analogue gravity in a number-conserving formalism
$\begin{split}\partial_{t}\theta_{c}+U\rho_{c}+\frac{1}{2}\left(\nabla\theta_{c}\right)\cdot\left(\nabla\theta_{c}\right)+f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)\\ +V({\bm{x}},t)=0.\end{split}$ (22)
@@
$f_{1}(\rho_{c},\theta_{c},{\bm{x}},t)$ $=$ $iU\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)},$
$f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)$ $=$ $2U\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+\frac{U}{2}\Big{(}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle e^{2i\theta_{c}}+\langle\hat{\chi}\hat{\chi}\rangle e^{-2i\theta_{c}}\Big{)}.$
which are related to real and complex parts of the “anomalous” density. We have omitted the “quantum pressure” term $-h(\rho_{c})=-\frac{1}{2}\rho_{c}^{-1/2}\left(\nabla^{2}\rho_{c}^{1/2}\right)$ on the left-hand side of the Bernoulli Eq. (22); we will generally neglect density gradients in the background(s) in what follows. Including these gradients would lead us to the concept of (nonlocal) rainbow metrics [48]. The indices 1 and 3 on $f_{1}$ and $f_{3}$ indicate their respective orders ${\cal O}(N^{0})$ and ${\cal O}(N^{-1})$; see also the note after Eq. (26) below.

### v3

@@ Abstract
We consider the quantum dynamics of Bose-Einstein condensates at absolute zero, and demonstrate that an analogue gravity model going beyond the standard linearized analogue gravity paradigm à la Unruh must take into account the backreaction of quasiparticle excitations onto the condensate background. This requires that one expands to second order in perturbation amplitude and thus takes the intrinsic nonlinearity of the theory into account. It is shown that, as a result, significant modifications of the standard paradigm occur. In particular, to obtain a fully Lorentz-covariant equation in curved spacetime for second-order perturbations, we demonstrate that it is necessary to introduce, to leading order in powers of the formal mean-field expansion parameter $N^{-1/2}$ (where $N$ is total particle number), a quantum-fluctuation-renormalized spacetime metric which substantially differs from the Unruh acoustic spacetime metric and, to subleading order $1/N$, two emergent vector fields and a mass term. Both the renormalized metric as well as the vector fields and the mass then keep track of the backreaction of the quasiparticles onto the condensate up to the order in powers of $N^{-1/2}$ considered. Finally, we apply our formalism to an analogue-cosmological Friedmann-Lemaître-Robertson-Walker metric and establish its renormalized form due to the quantum many-body backreaction exerted by the excitation cloud.
@@ I Introduction
It is of major importance in what follows that for a proper treatment of backreaction problems in quantum systems such as BECs, as was first pointed out in [30], it is necessary to maintain particle-number conservation, cf. Refs. [31, 32, 33, 34], which are obtaining variants of number-conserving formulations of Bogoliubov theory which go beyond the symmetry-breaking traditional Bogoliubov approach. In [30] the backreaction problem was defined by employing an expansion of the field operator in powers of $N^{1/2}$, where $N$ is the total number of particles (here either atoms or molecules) in the system. Then the backreaction of the quantum fluctuations can be elegantly treated as the subleading correction to the leading order in $N^{1/2}$ Gross-Pitaevskiǐ equation (GPE) [which is equivalent to a nonrelativistic variant of $\phi^{4}$ theory], incorporated by the modified coupled dynamical equation of the condensate and the noncondensate part. A particular lesson to be drawn from the analysis of [30] is that in a number-conserving description of the backreaction problem, the usual assumption of taking the expectation value of the quantized energy-momentum tensor, $\hat{T}_{\mu\nu}$, in a suitable quantum state, cf. Ref. [35], does not necessarily reproduce the correct, that is observable physical backreaction force.
@@
We will argue that the above procedure engenders a significant departure from the established formulation of the fluid perturbation equations as those of a scalar field in curved spacetime and that it is not much less straightforward to conclude that such a description can be envisaged if both backreaction and nonlinearity are taken into account. We will indeed demonstrate that for a Lorentz-invariant form of the nonlinear perturbation equation for a scalar field to exist, the latter will acquire, in subleading order, a mass term and two “emergent” vector fields. To leading order in the large $N$ expansion, the effective curved background metric acquires corrections stemming from the exchange of particles between condensate and excitation cloud. As our analysis in this paper is based on the Bogoliubov approximation, i.e., the number of particles forming the condensate must be much larger than that of forming the depleted cloud, the backreaction corrections to the standard GPE are by necessity small (for the Bogoliubov theory to apply the gas must be dilute), and consequently also the above mentioned differences from the standard linearized analogue gravity procedure (which we dub in what follows the Unruh paradigm), must be small. However, due to the high tunability of quantum gases, in particular with respect to the two-body coupling strength, by Feshbach resonances (cf., e.g., Refs. [43, 44]), one can potentially observe such deviations from the conventional analogue gravity paradigm in a highly controlled manner in cold quantum gas experiments.
The paper is organized as follows: In Section. II, we explain the formulation of the number-conserving approach we follow throughout the paper. Section III consists of the description of identifying the Lorentz invariant dynamical equation for the perturbed field. Section IV describes the quantum nonlinear corrections to the lowest order semiclassical background metric and Section V elaborates on these corrections for a Friedmann-Lemaître-Robertson-Walker metric. In Section VI we conclude, and an Appendix contains the explicit evaluation of the drain term driving condensate loss for a Bose gas in a hard-walled box trap which is quenched from zero to finite interaction coupling. We finally note that we use the term Unruh paradigm metric (or simply Unruh metric) instead of the more conventionally used “acoustic” or “sonic” spacetime metric, to clearly distinguish this linearized paradigm metric from the second-order metric we obtain below within our nonlinear approach to analogue gravity.
@@ II.1 Governing field equations
$i\partial_{t}\Psi=\left(-\frac{1}{2}\nabla^{2}+V+\lambda|\Psi|^{2}\right)\Psi,$ (1)
Here, $V$ is the external (trapping) potential, and $\lambda>0$ is the interaction coupling from $s$-wave scattering between the gas particles. The theory described by (1) is invariant under global U(1) transformations, which ensures the conservation of the total particle number $N$ according to the Noether theorem. The conservation law $\partial_{t}\rho+\nabla\cdot{\bm{J}}=0$ thus holds, where
@@
Here, the scaling behavior of each components taking the limit of very large $N$ while keeping $\lambda N=\mbox{constant}$, is given by [30]
@@
The above formal expansion (3) relies on the assumption that the number of noncondensed particles, $\delta N$, is much less than that of those forming the condensate part, and thus posits that $N_{0}\lesssim N$, also cf. Ref. [46] for the proof that in the mathematical limit $N\rightarrow\infty$ with $\lambda N$ held fixed, the GPE (5) below becomes exact. Here, $\chi$ is the fluctuation field associated with the noncondensed particles, and $\phi_{2}$ encompasses the correction of the condensate wave function due to the backreaction of the field $\chi$ onto the condensate. An expansion respecting the $N$ hierarchy is necessary to properly identify the fluctuation operator and the correction to the leading order Gross-Pitaevskiǐ orbital, see for details below.
@@
$i\frac{\partial}{\partial t}\phi_{0}=\left[-\frac{1}{2}\nabla^{2}+V+\lambda|\phi_{0}|^{2}\right]\phi_{0}.$ (5)
@@
$i\frac{\partial}{\partial t}\chi=\left(-\frac{1}{2}\nabla^{2}+V+2\lambda\rho_{0}\right)\chi+\lambda\phi_{\rm 0}^{2}\chi^{*},$ (6)
with $\rho_{0}=|\phi_{0}|^{2}$. The above equation and the corresponding Hermitian conjugate equation are the modified versions of the Bogoliubov–de Gennes equations in the non-number-conserving standard Bogoliubov approach. When we consider the next-order corrections to the condensate wave function due to the backreaction of the noncondensate part of the system, as was done in [33, 34], then it can be shown that the $\mathcal{O}(N^{0})$ corrections to the leading order GPE are identically zero [33]. However in the next order ${\mathcal{O}}(N^{-1/2})$, the standard GPE is modified due to the correction from the time-dependent backreaction of the noncondensate particles, as described by the fluctuation operator. To provide a description of the corresponding mechanism, we will follow [30, 36] and define a quantum-fluctuation-corrected condensate wave function
@@
$i\partial_{t}\phi_{c}=\Big{(}-\frac{1}{2}\nabla^{2}+V+\lambda|\phi_{\rm c}|^{2}+2\lambda|\chi|^{2}\Big{)}\phi_{c}+\lambda\chi^{2}\phi_{c}^{*}.$ (8)
This equation describes the coupled dynamics of condensed and noncondensed parts taking the dynamical corrections due to the backreaction into account. We note that the corrected order parameter $\phi_{c}$ has a dominant ${\cal O}(N^{1/2})$ term plus a subdominant correction of order $N^{-1/2}$, and for a proper formulation of backreaction, it is essential to take this correction into account. We pursue here the approach that backreaction is taken into account to this minimal order ${\cal O}(N^{-1/2})$ in the expansion (3). In principle, higher order corrections to the mean field $\phi_{c}$ in Eq. (7) in this formal expansion could be considered, and thus also higher order terms would occur in the effective fields of the analogue spacetime wave equation for second-order perturbations (37) which is derived below. We should finally also stress that we generally do not take further large $N$ approximations in the following expressions, above and beyond that of the field expansion order taken into account for defining $\phi_{c}$.
@@
$\langle\hat{\chi}({\bm{x}},t))\rangle\coloneqq 0\,\,\,\forall\,{\bm{x}},t\geq 0$ (10)
which constrains the space of states under consideration. The initial state for the Heisenberg picture evolution is here chosen as the noninteracting $\lambda=0$ state which has zero depletion $\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle=0\,\,(t=0)$ and $\phi_{2}=0$ (zero backreaction) [36], and can be set up in experiment in a well controlled manner by tuning the system to the zero-crossing of a Feshbach resonance. At $t=0$, we quench to a finite value of $\lambda$ and follow the subsequent evolution, maintaining the condition (10) corresponding initially to the noninteracting vacuum.
The quantization choice to regard $\hat{\chi}$ as the fundamental bosonic quantum field instead of $\hat{\Psi}$, does not maintain number conservation on the level of the full field operator of the bosonic particles: Due to $\langle\hat{\Psi}\rangle=\phi_{c}$, where however now $\phi_{c}$ is the backreaction-renormalized condensate mode in Eq. (7), the U(1) symmetry is still broken on the quantum level, as in the original Bogoliubov theory [27]. This choice, to promote $\chi$ to be the “fundamental” quantum field, however yields a set of equations amenable to a concrete solution the backreaction problem  [^2] We note in this connection that we were unable to locate an explicit backreaction calculation following the formally rigorous and in this regard very appealing approach of Ref. [33], which maintains $\langle\hat{\Psi}\rangle=0$ for any fixed $N$ state.. Furthermore, and most importantly in our context, it has been shown in [36] that with this vacuum choice for the quantized field $\hat{\chi}$, number conservation for condensate plus excitations is maintained, see Eqs. (18) and (19) below, and up to the order in $N^{-1/2}$ considered.
@@
$i\partial_{t}\phi_{c}=\left[-\frac{1}{2}\nabla^{2}+V+\lambda|\phi_{\rm c}|^{2}+2\lambda\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle\right]\phi_{c}+\lambda\langle\hat{\chi}^{2}\rangle\phi_{c}^{*}.$ (11)
@@ II.2 Quantum backreaction
$\rho$ $=$ $|\phi_{c}|^{2}+\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+{\mathcal{O}}(N^{-1/2}),$ (14)
${\bm{J}}$ $=$ $\frac{1}{2i}\Big{(}\phi^{*}_{c}\nabla\phi_{c}-\phi_{c}\nabla\phi^{*}_{c}\Big{)}$ (15)
@@
We then have $\rho=\rho_{c}+\rho_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$ and ${\bm{J}}={\bm{J}}_{c}+{\bm{J}}_{\hat{\chi}}+{\mathcal{O}}(N^{-1/2})$, with ${\bm{J}}_{c}=\frac{1}{2i}(\phi^{*}_{c}\nabla\phi_{c}-\phi_{c}\nabla\phi^{*}_{c})$.
Neither condensate nor depletion are separately conserved. In particular, we have for condensate and depletion cloud, respectively [36, 37]
$\partial_{t}\rho_{c}+\nabla\cdot{\bm{J}}_{c}$ $=$ $i\lambda\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}$ (18)
$\partial_{t}\rho_{\hat{\chi}}+\nabla\cdot{\bm{J}}_{\hat{\chi}}$ $=$ $-i\lambda\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)}.$ (19)
We conclude that the theory is indeed conserving up to order ${\mathcal{O}}(N^{0})$ for density and current. This defines the extent to which our approach is number-conserving. We observe that this conservation order directly corresponds to the order ${\cal O}(N^{-1/2})$ for the mean-field correction $\phi_{2}$ in Eq. (7), and is not derived by any further approximation. The number of particles in condensate and in the depletion cloud of quantum excitations are not separately conserved but only their respective sums, and thus the total density and current. The separate conservation equations for condensate and excitations acquire drain (source) terms which stem from backreaction. We furthermore emphasize that number conservation holds here for the averaged quantities density and current. Fluctuations of leading relative order ${\cal O}(N^{-1/2})$ persist, as we break number conservation on the microscopic quantum level of the elementary bosons by imposing $\langle\hat{\Psi}\rangle=\phi_{c}$, which corresponds to a coherent state for large $N$ with Poissonian statistics. Note also that the prescription of quantization we employ in Eqs. (9) implies that $\hat{\Psi}$ and $\hat{\Psi}^{\dagger}$ do not satisfy the canonical bosonic commutators. The role of the quantized bosonic fields is taken over by $\hat{\chi},\hat{\chi}^{\dagger}$. This is akin to what happens in the original Bogoliubov approach [27].
@@ III Analogue gravity in a number-conserving formalism
$\begin{split}\partial_{t}\theta_{c}+\lambda\rho_{c}+\frac{1}{2}\left(\nabla\theta_{c}\right)^{2}+f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)+V({\bm{x}},t)=0.\end{split}$ (22)
@@
$\begin{split}f_{1}(\rho_{c},\theta_{c},{\bm{x}},t)=i\lambda\Big{(}\phi^{2}_{c}\langle\hat{\chi}^{\dagger}\hat{\chi}^{\dagger}\rangle-\phi^{*2}_{c}\langle\hat{\chi}\hat{\chi}\rangle\Big{)},\\ f_{3}(\rho_{c},\theta_{c},{\bm{x}},t)=2\lambda\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle+\frac{\lambda}{2}\Big{(}\langle\hat{\chi}\hat{\chi}\rangle e^{-2i\theta_{c}}+{\rm h.c.}\Big{)}.\end{split}$ (23)
which depend on real and imaginary parts of the “anomalous” excitation density $\langle\hat{\chi}\hat{\chi}\rangle$. We have omitted the “quantum pressure” term $-h(\rho_{c})=-\frac{1}{2}\rho_{c}^{-1/2}\left(\nabla^{2}\rho_{c}^{1/2}\right)$ on the left-hand side of the Bernoulli Eq. (22); we will generally neglect density gradients in the background(s) in what follows. Including these gradients would lead us to the concept of (nonlocal) rainbow metrics [48]. The indices 1 and 3 on $f_{1}$ and $f_{3}$ indicate their respective orders ${\cal O}(N^{0})$ and ${\cal O}(N^{-1})$; see also the note after Eq. (26) below.